Find the total admissions for the top-grossing movie if they were 11/24 of the total admissions for the top grossing movie of all time, which had a total admission of 205 million.
Answer by stanbon(75887)

Find the total admissions for the top-grossing movie if they were 11/24 of the total admissions for the top grossing movie of all time, which had a total admission of 205 million.
-----------------------
Ans: (11/24)(205 million) = 0.4583*205 million = 93.96 million
================
Cheers,
Stan H.